OpenStudy (anonymous):

Could I find two missing angle measures if I know some of the side lengths of a right triangle?

OpenStudy (anonymous):

AB = 8 BC = 6 CA = 10 Angle B = 90 degrees

OpenStudy (anonymous):

Can someone help show the steps to find the missing angles? like step by step?

OpenStudy (anonymous):

@Preetha @chmvijay @abb0t

OpenStudy (chmvijay):

@goformit100 @satellite73 @skullpatrol @.Sam. @amistre64 @uri @kelliegirl33 @SWAG can u help him out

OpenStudy (anonymous):

Thanks @chmvijay :)

OpenStudy (anonymous):

@thomaster Do you have any idea? :/

OpenStudy (anonymous):

Sure use trig

OpenStudy (anonymous):

Are you familiar with SOH CAH TOA?

OpenStudy (anonymous):

a tad, but i never fully understood it

OpenStudy (anonymous):

Ok look SOH CAH TOA is a memory trick to remember the following...

OpenStudy (anonymous):

\[ \sin A = \frac{ opposite leg }{ hypotenuse }\] \[ \cos A = \frac{ adjacent leg }{ hypotenuse }\] \[ \tan A = \frac{ opposite leg }{ adjacent leg }\]

OpenStudy (anonymous):

A is the angle. Opposite leg means the leg across from the angle. Adjacent leg means the leg that is "hooked into" or touchig the angle (the leg not the hypotenuse) and hypotentuse the longest side (the one across from the right angle)/.

OpenStudy (anonymous):

SO if you pick an angle that you want to find. Which one do you want to start with?

OpenStudy (anonymous):

lets try angle a

OpenStudy (anonymous):

@Paynesdad

OpenStudy (anonymous):

Ok so now since you are looking for the angle you need to know at least two sides do in fact you know all three so let's pick any two of the three so which ones do you want to pick?

OpenStudy (anonymous):

ab and bc?

OpenStudy (anonymous):

@Paynesdad

OpenStudy (anonymous):

So compared to A is AB the opp side, adj side or the hyp?

OpenStudy (anonymous):

adj side right?

OpenStudy (anonymous):

right and what about BC?

OpenStudy (anonymous):

compared to a? opp side?

OpenStudy (anonymous):

@Paynesdad

OpenStudy (anonymous):

Sorry...Had to take care of something...yep BC is opp of A so you have the adjacent side and the opp side which trig function doies that indicate that you should use?

OpenStudy (anonymous):

tan a im guessing

OpenStudy (anonymous):

@Paynesdad

OpenStudy (anonymous):

@Luigi0210 do you have any idea whats next?

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

so you write....Tan of the angle equals the opp leg value over the adjacent leg value...or Tan A = 6/8

OpenStudy (anonymous):

would that change to 3/4?

OpenStudy (anonymous):

Well yes but I would just write it as a decimal because you will use you calculator soon anyway.

OpenStudy (anonymous):

so Tan A = 0.75

OpenStudy (anonymous):

Now you have to remove the Tan from the A or as I like to say UNDO the tangent. SO in algebra if you have 6x = 12 and you want to get rid of the 6 times x so it is just x ... you divide both sides by 6 to undo the 6 times ... right?

OpenStudy (anonymous):

yes, so divide both sides by tan?

OpenStudy (anonymous):

no because tan isn't a number it is a function Tan A doesn't mean Tan times A ... it is a special function of its own... Tan A means the Tangent of A. so to undo Tan A, I have to have a reverse function of Tan just like the reverse function of times is divide and vice versa

OpenStudy (anonymous):

so the reverse function or the undo function of"Tangent of" is"Inverse Tangent of" or in symbols the inverse of Tan is \[tan^-1\]

OpenStudy (anonymous):

So you want to undo Tan in the equation Tan A = 0.75 so ... you take the Tan^-1 of both sides like this... \[tan^-1 (tan A) = tan^-1 (0.75) \]

OpenStudy (anonymous):

Now on the left hand side the tan-1 and the tan cancel just like in 6x/6 the 6 times and the 6 divide cancel to leave just x. so the left hand side of the equation above goes from \[tan^-1 (tan A) =\] to just A =

OpenStudy (anonymous):

and on the right hand side you have \[tan^-1 (0.75)\] which you just put in your calculator to get a value ... make sure your calculator is in degrees mode. The value you get in your calculator should be the size of A in degrees. so you final answer is A = (whatever number is in your calculator) it should be bigger than 30degrees but smaller than 45.

OpenStudy (anonymous):

Alright, well i know what to do, but i just dont have a scientific calculator, i have a plain one:(

OpenStudy (anonymous):

@Paynesdad

OpenStudy (anonymous):

Do you have a calc on your computer?

OpenStudy (anonymous):

yes but its a simple one, only basic functions and the "M+" and "MC", and "M-"

OpenStudy (anonymous):

Usually in the view menu of that calculator is a choice to change it over to a scientific calc. you could try that.

OpenStudy (anonymous):

oh sweet okay it worked, so what would i plug in exactly?

OpenStudy (anonymous):

Make sure the calc is set for degrees.

OpenStudy (anonymous):

so you should have shown the following work...

OpenStudy (anonymous):

deg means degrees right? thats what it says

OpenStudy (anonymous):

\[tan A = \frac{6}{8} = 0.75\] \[tan^-1 (tan A) = tan^-1(0.75)\] \[A = ______\]

OpenStudy (anonymous):

yes deg means geree mode...Now type in your calc 0.75

OpenStudy (anonymous):

then the "Inv" (Inverse) button.

OpenStudy (anonymous):

then "Tan"

OpenStudy (anonymous):

you could then hit equal... What do you get?

OpenStudy (anonymous):

none of the calculators have inv on it:/

OpenStudy (anonymous):

@Paynesdad

OpenStudy (anonymous):

@Luigi0210 im on the last part of this, and he left. you know the rest?

OpenStudy (anonymous):

@dan815 hi luigi said to get you to help with the end of this problem?:)

OpenStudy (dan815):

is it the problem at the start of this conversation?

OpenStudy (anonymous):

Sorry I am back... not sure why the calculator does not work. Use the google calc that you have but type inv then tan then 0.75... try that

OpenStudy (anonymous):

okay i got 0.64, would that be correct?

OpenStudy (anonymous):

That's in rad click the rad button at the top and do it again in degrees.

OpenStudy (anonymous):

36.86?

OpenStudy (anonymous):

There you go...Now that you have two angles and you know all three add to 1800 you should be able to find the third.

OpenStudy (anonymous):

54?

OpenStudy (anonymous):

yep! good job... Glad we could finally get this done.

OpenStudy (anonymous):

thank you so much you dont understand how much that helped me!